OpenStudy (anonymous):

If the radius of a tennis ball is 3.5 cm, what percent of the can is occupied by air, not including the air inside of the balls?

OpenStudy (agent0smith):

This question is incomplete, post the whole question. But you will probably need: Volume of a sphere: \[\Large V _{sphere} = \frac{ 4 }{ 3 } \pi r^3\] Volume of a cylinder \[\Large V _{cylinder} = \pi r^2 h\] You'll have to subtract the volumes of the tennis balls from the volume of the cylinder, then to get it as a percentage: divide that amount by the volume of the cylinder and times by 100. to find "what percent of the can is occupied by air, not including the air inside of the balls?"

OpenStudy (anonymous):

The entire question follows: A standard tennis ball can is a cylinder that hold 3 tennis balls. If the radius of a tennis ball is 3.5 cm, what percent od the can is occupied by air, not including the air inside the balls?

OpenStudy (anonymous):

of*

OpenStudy (agent0smith):

So you just need to do exactly as I said above, so subtract the volume of the three balls from the volume of the cylinder. Note that the cylinder will have a height of 6 times the radius of the balls; see if you can figure out why (draw a diagram for yourself)...